OpenStudy (lilai3):

what is x to the 7th power times x to the negative 3 power

OpenStudy (anonymous):

\[x^7*x^{-3}\]

OpenStudy (lilai3):

iz da answer one over x to the 14th power?

OpenStudy (anonymous):

\[x^{-a}= \frac{1}{x^a}\]

OpenStudy (lilai3):

what is x to the negave 7th power times x to the third power?

OpenStudy (lilai3):

oh @Romero, is that the answer?

OpenStudy (anonymous):

No I'm telling you how you can rewrite negative powers ....

OpenStudy (lilai3):

oh pish uhm... one to the x to the 4th power?

OpenStudy (anonymous):

\[x^7*x^{-3}=\frac{x^7}{x^3}\]

OpenStudy (lilai3):

oh iz dat the answer my bad....

OpenStudy (anonymous):

when ever you have a negative power....switch it back to a positive by bringing it down as a fraction

OpenStudy (lilai3):

wait is that the first prmblm or second

OpenStudy (anonymous):

Can you post another question...im really confused at what question you need help on

OpenStudy (lilai3):

is that top answerx7∗x−3=x7x3 is it the question i poseted?

OpenStudy (anonymous):

I really don't understand what you are saying....you are not using the appropriate signs

OpenStudy (lilai3):

oh i just copied and pasted ....-__-